# Hidden Bose-Einstein Singularities in Correlated Electron Systems

### v1

@@ Appendix A Numerical Procedures
Numerical calculations of Eqs. (11) and (18) have been performed iteratively by choosing the noninteracting and mean-field Green’s functions as the starting points, respectively. Specifically for the normal state, we divide Green’s function in $\tau$ space with $\tau\in[0,\beta]$ as
@@
$G_{0}(k,\tau)=-\frac{e^{-\tau(k^{2}-\mu_{0})}}{e^{-\beta(k^{2}-\mu_{0})}+1},$ (22)
and $\delta G(k,\tau)$ denotes the correlation part updated iteratively starting from $\delta G(k,\tau)\!=\!0$. Variable $k$ is expressed as
$k=1+\sinh x,$ (23a)
and discretized with an equal interval in $x$ over $100\!\leq\!k\!\leq\!k_{\rm c}\!=\!$ 10; the number of points are $N_{k}\!=\!50\!\sim\!200$. Similarly, $\tau$ is expressed as
$\tau=\frac{\beta}{2}(1+\tanh x),$ (23b)
and discretized with an equal interval in $x$ over $\tau_{\rm c}\!\leq\!\tau\!\leq\!\beta\!-\!\tau_{\rm c}$ with $\tau_{\rm c}\!=\!10^{-8}$; the number of points are $N_{\tau}\!=\!50\!\sim\!400$.
The correlation functions of Eq. (4) are calculated in $\tau$ space based on the formula,
$\chi_{GG}(q,\tau)=\frac{1}{V}\sum_{\bf k}G(k,\tau)G(|{\bf k}-{\bf q}|,\beta-\tau),$ (24a)
$\chi_{G\bar{G}}(q,\tau)=-\frac{1}{V}\sum_{\bf k}G(k,\tau)G(|{\bf k}-{\bf q}|,\tau).$ (24b)
The angular integrals of $s\equiv{\bf k}\!\cdot\!{\bf q}/kq$ in Eq. (24) are performed over
${\max}\!\left(-1,\frac{k^{2}\!+\!q^{2}\!-\!k_{\rm c}^{2}}{2kq}\right)\!\leq s\leq 1,$ (25)
by dividing the interval at $(k^{2}\!+\!q^{2}\!-\!1)/2kq$ when it lies in it. The integration over each interval $s_{0}\!\leq\!s\!\leq s_{1}$ is performed by expressing
$s=\frac{s_{0}\!+\!s_{1}}{2}+\frac{s_{0}\!-\!s_{1}}{2}\tanh(\sinh x),$ (26)
and discretizing $x$ at an equal interval over $-2.5\!\leq\!x\!\leq\!2.5$; the number of integration points are $N_{s}\!=\!20\!\sim\!30$. The radial integrals of ${\bf k}$ in Eq. (24) are performed similarly over
${\rm max}(0,q-k_{\rm c})\leq k\leq k_{\rm c},$ (27)
by dividing the interval at the singular points $|1\!-\!q|$, $1$, $1\!+q$ when they lie in it, and making a change of variables similar to Eq. (26) for each interval.
Variable $q$ of Eq. (24) runs over $[0,2k_{\rm c}]$, which is divided at $q\!=\!2,k_{\rm c}$ into three regions. Each region is then discretized by the formula (26) with $s\rightarrow q$ with the relative number of points $4:3:3$; the total number of points is $N_{q}\!=\!50\!\sim\!200$.
The functions of Eq. (24) are then transformed into the $\omega_{\ell}$ space based on the method developed by Haussmann et al.HRCZ07 To be specific, we express any function $\chi(\tau)$ with $\tau\!\in\![\tau_{j},\tau_{j+1})$ by the cubic splineNRC
@@
and carry out the Fourier transform by
@@
The boson Matsubara frequencies $\omega_{\ell}\!=\!2\pi\ell T$ ($\ell\!=\!0,1,2,\cdots$) for $\ell\!>\!\ell_{\rm c}\!\sim\!10$ are chosen so that they are described approximately by $\omega_{\ell}\!\approx\!\sinh x_{j}$ in terms of equally spaced points $x_{j}$ ($j\!=\!\ell_{\rm c}\!+\!2,\cdots,N_{\tau}$) up to $\omega_{\ell}\!\approx\!10^{7}$; the number of points are $N_{\omega}\!=\!50\sim 200$. Function $\bar{\chi}(\omega)$ with $\omega\!\in\![\omega_{\ell_{j}},\omega_{\ell_{j+1}})$ is then approximated by the cubic spline as
@@
The accuracy of Eq. (29) can be checked by the inverse transform
@@
and replace the upper limit $\infty$ in Eq. (33b) by $k_{\rm c}\!\sim\!10$ to obtain enough accuracy. The sum of Eq. (34) for $\nu\!\lesssim\!10$ can be evaluated by using Faulhaber’s formula. Equation (32) after the double Fourier transforms thereby obtained has been confirmed to reproduce Eq. (28) excellently.
Functions $\chi_{GG}(\vec{q}\,)$ and $\chi_{G\bar{G}}(\vec{q}\,)$ calculated from Eq. (24) by Eq. (29) are then substituted in Eq. (5). The sum over $\ell$ is performed in $\tau$ space by using the Fourier transforms of
@@
$\Sigma_{\rm c}(k,\tau)=$ $\,\frac{1}{V}\sum_{{\bf q}}\bigl{[}U_{\rm ph}(q,\tau)G(|{\bf k}-{\bf q}|,\tau)$
$\,-U_{\rm pp}(q,\tau)G(|{\bf k}-{\bf q}|,\beta-\tau)\bigr{]},$ (36)
where the subscript of $\Sigma_{\rm c}$ denotes the correlation part. On the other hand, the angular and radial integrals of ${\bf q}$ in Eq. (36) are carried out in the same way as those over ${\bf k}$ in Eq. (24). Possible singular points in the radial integral over $q\!\in\![0,k\!+\!k_{\rm c}]$ are $|1\!-\!k|$, $1\!+\!k$, and $k_{\rm c}\!-\!k$.
Equation (36) thereby obtained is transformed into the $\varepsilon_{n}$ space in the same way as described from Eq. (32) through Eq. (34). The fermion Matsubara frequencies $\varepsilon_{n}\!=\!(2n\!+\!1)\pi T$ ($n\!=\!0,1,2,\cdots$) for $n\!>\!n_{\rm c}\!\sim\!10$ are chosen so that they are described approximately by $\varepsilon_{n}\!\approx\!\sinh x_{j}$ in terms of equally spaced points $x_{j}$ ($j\!=\!n_{\rm c}\!+\!2,\cdots,N_{\tau}$) up to $\varepsilon_{n}\!\approx\!10^{7}$; the number of points are $N_{\varepsilon}\!=\!50\sim 200$. Self-energy $\Sigma(\vec{k})$ thereby obtained is used to calculate $\delta G(\vec{k})\!\equiv\!G(\vec{k})\!-\!G_{0}(\vec{k})$, i.e., the correlation part of Green’s function. This $\delta G(\vec{k})$ is then transformed into the $\tau$ space by the procedure from Eq. (28) to Eq. (30) to obtain $\delta G(k,\tau)$ of the next iteration.
Below $T\!=\!T_{1{\rm PR}}$, we need to incorporate the second term of Eq. (10) to obtain any finite result. Calculations have been performed for several trial $\lambda_{\rm n}$’s to identify the value of $\lambda_{\rm n}$ at which $U\chi_{G\bar{G}}(\vec{0})\!=\!1$ is met. It has turned out numerically that the value of $U\chi_{G\bar{G}}(\vec{0})$ changes almost linearly in terms of $\lambda_{\rm n}$ around $U\chi_{G\bar{G}}(\vec{0})\!=\!1$, so that the desired value of $\lambda_{\rm n}$ has been identified with no fundamental difficulties.
Similar calculations have been performed for the superconducting state iteratively starting from the mean-field solutions. The convergences are not bad at low temperatures of $T\!\lesssim\!0.2T_{{\rm c}0}$ even for the calculations of using Eq. (17). As we approach $T_{\rm c}\!\approx\!0.0624$, however, they become more and more unstable. Hence, we have set $\lambda_{\bar{G}}\!=\!\tilde{\lambda}_{F}\!>\!0$ and $\lambda_{G}\!=\!\lambda_{F}\!=\!\tilde{\lambda}_{G}\!=\!0$ from the beginning based on the results at low temperatures. This has enabled us to obtain the convergence near $T_{\rm c}$, which however has been reached only after many iterations.

### v2

@@ Appendix B Numerical Procedures
Numerical calculations of Eqs. (11) and (18) have been performed iteratively by choosing the noninteracting and mean-field Green’s functions as the starting points, respectively. Specifically for the normal state, we divide Green’s function in the $\tau$ space with $\tau\in[0,\beta]$ as
@@
$G_{0}(k,\tau)=-\frac{e^{-\tau(k^{2}-\mu_{0})}}{e^{-\beta(k^{2}-\mu_{0})}+1}$ (28)
and $\delta G(k,\tau)$ denotes the correlation part updated iteratively starting from $\delta G(k,\tau)\!=\!0$. The variable $k$ is expressed as
$k=1+\sinh x$ (29a)
and discretized with an equal interval in $x$ over $100\!\leq\!k\!\leq\!k_{\rm c}\!=\!$ 10; the number of points is $N_{k}\!=\!50$-$200$. Similarly, $\tau$ is expressed as
$\tau=\frac{\beta}{2}(1+\tanh x)$ (29b)
and discretized with an equal interval in $x$ over $\tau_{\rm c}\!\leq\!\tau\!\leq\!\beta\!-\!\tau_{\rm c}$ with $\tau_{\rm c}\!=\!10^{-8}$; the number of points is $N_{\tau}\!=\!50$-$400$.
The correlation functions of Eq. (4) are calculated in the $\tau$ space on the basis of the equations
$\chi_{GG}(q,\tau)=$ $\,\frac{1}{V}\sum_{\bm{k}}G(k,\tau)G(|{\bm{k}}-{\bm{q}}|,\beta-\tau),$ (30a)
$\chi_{G\bar{G}}(q,\tau)=$ $\,-\frac{1}{V}\sum_{\bm{k}}G(k,\tau)G(|{\bm{k}}-{\bm{q}}|,\tau).$ (30b)
The angular integrals of $s\equiv{\bm{k}}\!\cdot\!{\bm{q}}/kq$ in Eq. (30) are performed over
${\max}\!\left(-1,\frac{k^{2}\!+\!q^{2}\!-\!k_{\rm c}^{2}}{2kq}\right)\!\leq s\leq 1$ (31)
by dividing the interval at $(k^{2}\!+\!q^{2}\!-\!1)/2kq$ when the value lies in the interval. The integration over each interval $s_{0}\!\leq\!s\!\leq s_{1}$ is performed by expressing
$s=\frac{s_{0}\!+\!s_{1}}{2}+\frac{s_{0}\!-\!s_{1}}{2}\tanh(\sinh x)$ (32)
and discretizing $x$ at an equal interval over $-2.5\!\leq\!x\!\leq\!2.5$; the number of integration points is $N_{s}\!=\!20$-$30$. The radial integrals of ${\bm{k}}$ in Eq. (30) are performed similarly over
${\rm max}(0,q-k_{\rm c})\leq k\leq k_{\rm c}$ (33)
by dividing the interval at the singular points $|1\!-\!q|$, $1$, $1\!+q$ when they lie in the interval, and changing variables similarly to Eq. (32) for each interval.
The variable $q$ of Eq. (30) runs over $[0,2k_{\rm c}]$, which is divided at $q\!=\!2,k_{\rm c}$ into three regions. Each region is then discretized by Eq. (32) with $s\rightarrow q$ with the relative number of points $4:3:3$; the total number of points is $N_{q}\!=\!50$-$200$.
The functions of Eq. (30) are then transformed into the $\omega_{\ell}$ space on the basis of the method developed by Haussmann et al.HRCZ07 To be specific, we express any function $\chi(\tau)$ with $\tau\!\in\![\tau_{j},\tau_{j+1})$ through the cubic spline,NRC
@@
and carry out the Fourier transform as
@@
The boson Matsubara frequencies $\omega_{\ell}\!=\!2\pi\ell T$ ($\ell\!=\!0,1,2,\cdots$) for $\ell\!>\!\ell_{\rm c}\!\sim\!10$ are chosen so that they are described approximately by $\omega_{\ell}\!\approx\!\sinh x_{j}$ in terms of equally spaced points $x_{j}$ ($j\!=\!\ell_{\rm c}\!+\!2,\cdots,N_{\tau}$) up to $\omega_{\ell}\!\approx\!10^{7}$; the number of points is $N_{\omega}\!=\!50$-$200$. The function $\bar{\chi}(\omega)$ with $\omega\!\in\![\omega_{\ell_{j}},\omega_{\ell_{j+1}})$ is then approximated by the cubic spline as
@@
The accuracy of Eq. (35) can be examined by the inverse transform
@@
and replace the upper limit $\infty$ in Eq. (39b) by $k_{\rm c}\!\sim\!10$ to obtain sufficient accuracy. The sum of Eq. (40) for $\nu\!\lesssim\!10$ can be evaluated using Faulhaber’s formula. Equation (38) after the double Fourier transforms thereby obtained has been confirmed to reproduce Eq. (34) excellently.
The functions $\chi_{GG}(\vec{q}\,)$ and $\chi_{G\bar{G}}(\vec{q}\,)$ calculated from Eq. (30) using Eq. (35) are then substituted in Eq. (5). The sum over $\ell$ is calculated in the $\tau$ space by the Fourier transforms of
@@
$\Sigma_{\rm c}(k,\tau)=$ $\,\frac{1}{V}\sum_{{\bm{q}}}\bigl{[}U_{\rm ph}(q,\tau)G(|{\bm{k}}-{\bm{q}}|,\tau)$
$\,-U_{\rm pp}(q,\tau)G(|{\bm{k}}-{\bm{q}}|,\beta-\tau)\bigr{]},$ (42)
where the subscript of $\Sigma_{\rm c}$ denotes the correlation part. On the other hand, the angular and radial integrals of ${\bm{q}}$ in Eq. (42) are carried out in the same way as those of ${\bm{k}}$ in Eq. (30). Possible singular points in the radial integral of $q\!\in\![0,k\!+\!k_{\rm c}]$ are $|1\!-\!k|$, $1\!+\!k$, and $k_{\rm c}\!-\!k$.
Equation (42) thereby obtained is transformed into the $\varepsilon_{n}$ space in the same way as described for Eqs. (38)-(40). The fermion Matsubara frequencies $\varepsilon_{n}\!=\!(2n\!+\!1)\pi T$ ($n\!=\!0,1,2,\cdots$) for $n\!>\!n_{\rm c}\!\sim\!10$ are chosen so that they are described approximately by $\varepsilon_{n}\!\approx\!\sinh x_{j}$ in terms of equally spaced points $x_{j}$ ($j\!=\!n_{\rm c}\!+\!2,\cdots,N_{\tau}$) up to $\varepsilon_{n}\!\approx\!10^{7}$; the number of points is $N_{\varepsilon}\!=\!50$-$200$. The self-energy $\Sigma(\vec{k})$ thereby obtained is used to calculate $\delta G(\vec{k})\!\equiv\!G(\vec{k})\!-\!G_{0}(\vec{k})$, i.e., the correlation part of Green’s function. This $\delta G(\vec{k})$ is then transformed into the $\tau$ space by the procedure for Eqs. (34)-(36) to obtain $\delta G(k,\tau)$ of the next iteration.
Below $T\!=\!T_{1{\rm PR}}$, we need to incorporate the second term of Eq. (10) to obtain any finite result. Calculations have been performed for several trial $\lambda_{\rm n}$’s to identify the value of $\lambda_{\rm n}$ at which $U\chi_{G\bar{G}}(\vec{0})\!=\!1$ is met. It has turned out numerically that the value of $U\chi_{G\bar{G}}(\vec{0})$ changes almost linearly with $\lambda_{\rm n}$ around $U\chi_{G\bar{G}}(\vec{0})\!=\!1$, so that the desired value of $\lambda_{\rm n}$ has been identified with no fundamental difficulties.
Similar calculations have been performed for the superconducting state iteratively starting from the mean-field solutions. The convergences are good at low temperatures of $T\!\lesssim\!0.2T_{{\rm c}0}$ even for the calculations using Eq. (17). As we approach $T_{\rm c}\!\approx\!0.0624$, however, the iterations become increasingly unstable. Hence, we have set $\lambda_{\bar{G}}\!=\!\tilde{\lambda}_{F}\!>\!0$ and $\lambda_{G}\!=\!\lambda_{F}\!=\!\tilde{\lambda}_{G}\!=\!0$ from the beginning on the basis of the results obtained at low temperatures. This has enabled us to obtain the convergence near $T_{\rm c}$, which however has been reached only after many iterations.